What is the equilibrium concentration of chloroacetic acid, HC2H2O2Cl, in a solution prepared by dissolving 0.0124 mol of HC2H2O2Cl in 1.50 L of water? For chloroacetic acid, Ka = 1.4
Answer #1

eq conc of chloracetic acid =5.49*10^-3   M

explanation= Ka= 1.4*10^-3 = (A^2)/((0.0124/1.5)-A)           So A= 2.77*10^(-3)

concn at equalibrium= ( .0124/1.5) - A = 5.49*10^(-3) M